

********************************************************************
UARS WEEKLY STATUS REPORT - 11 FEB 2005
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
64% of each day's data amounting to about 444 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for time periods with UARS telemetry or tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004269 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004270 and 2004279 are currently waiting archival at GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004279. MEPS data between 2004280 and 2004289
are currently being processed.
PEM continues to receive NMC data files from NOAA. NMC host machine
changed during mid-January which caused some NMC data transfer failures;
however, this situation was corrected and recovery completed with the
assistance of Roger Lin at NOAA. No problems have been encountered since
updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 29 January 2005 (Orbit 73255, GMT Day 029) through
04 February 2005 (Orbit 73361, GMT Day 035).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 030/2336 73285 POWER OFF
SUSIM 034/2039 73344 SUSIM RESET
The beta angle ranged between the angles of 38.7 degrees to 28.1
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 1.0 degrees on 12 February 2005 (DOY 043).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.65 to 1.00 degrees C. Additional
battery performance data for each battery on 28 January 2005 and 04
February 2005 are provided in Appendix D.
The clock error ranged from 5.0 to -6.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 01
PEM = 01
SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 06
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week batteries have performed nominally and the MPS continues to
utilize the VT-7 charge rate.
The FOT has completed their CMS software acceptance testing. The next
step in the process is for the official release of the software to be
installed on the backup CMS. This will be done by the software
development team next week. Once that is accomplished, the FOT will
run a few simulated days of operations to confirm the install. If all
goes well, then the new release will be put onto the operational CMS.
MEETINGS
The FOT attended the UARS monthly status review on 01 February 2005 to
discuss the significant events for the month of January.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 12 February 2005 (DOY 043).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 04 February 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
73285 030/2336 05-033 PEM POWER OFF
73336 034/0841 05-034 OBC ATC NOOP:LOC=705
73344 034/2039 05-035 SUSIM SUSIM RESET
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 04 February 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 04 February 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period.
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 04 February 2005
28 January 2005 - GMT Day 028 Beta = 38.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.79/-2.73 +00.0/+00.0 06.9 0.00 0.0
2 156.8/ 22.4 +9.73/+8.40 +27.2/-14.8 26.9 1.08 19.1
3 22.4/ -78.4 +7.41/+6.44 +25.6/-14.4 26.7 1.07 18.2
04 February 2005 - GMT Day 035 Beta = 28.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.48/-2.42 +00.0/+00.0 07.0 0.00 0.0
2 -5.6/ -67.2 +10.06/+9.06 +28.4/-15.2 26.7 1.07 20.9
3 22.4/ -72.8 +7.74/+7.09 +26.8/-14.4 26.7 1.05 20.1
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 04 February 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
5-005 FOT 73302 032/0228 30832 Generic Late Acquisition
05-006 WSGT 73303 032/0336 38406 WS not sending UPDs for multip
05-005 FOT 73317 033/0118 30832 Generic Late Acquisition
05-005 FOT 73341 034/1639 30832 Generic Late Acquisition
05-007 FOT 73344 034/2039 N/A SUSIM RESET
05-005 FOT 73345 034/2256 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-001 FOT See List Jan 05 30832 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-005 FOT See List Feb 05 30832 Generic Late Acquisition
05-002 MPS 73008 012/1251 NA BATT 3 Diff Voltage High
05-006 WSGT 73303 032/0336 38406 WS not sending UPDs for multip
05-007 FOT 73344 034/2039 N/A SUSIM RESET